OpenStudy (anonymous):

If the slope of a line is 1/2 and the y-intercept is 6, what is the x-intercept of the same line?

OpenStudy (anonymous):

y=1/2x+6 .... i don't know how to get started

OpenStudy (texaschic101):

to find the x intercept, sub in 0 for y y = 1/2x + 6 0 = 1/2x + 6 -1/2x = 6 x = 6/(-1/2) x = 6 * -2 x = -12 so your x intercept is -12...or (-12,0)

OpenStudy (anonymous):

thanks! why did i forget about that? -.-

OpenStudy (texaschic101):

anytime :) glad to help